[6] (5 points) There exist four Boolean functions of a single variable and 16 functions of two variables How many functions oTHIS IS ONE QUESTION
Answer #1

A single variable has 4 boolean functions and 2 two variable has 16 functions.

Now,

The no. of boolean function with 5 variable are = 22^n

= 22^5 (where, n is the no. of variable, i.e. n = 5)

=232

= 4294967296

The no. of boolean functions wit n variables are, = 22^n